Fred measures (x, y, z, t) ⇒ Jane measures (x′, y′, z′, t′) x, y, z, t and relative speed v along x axis x′, y′, z′, t′ and relative speed − v along x axis

\begin{aligned}x′ & = x − vt \\ y′ & = y \\ z′ & = z \\ t′ & = t \end{aligned}